OpenStudy (anonymous):

\[x-\sqrt[4]{x} \div x-1\]

OpenStudy (anonymous):

need to simplify

OpenStudy (anonymous):

\[\frac{x-x^{1/4}}{x-1}\]

OpenStudy (anonymous):

yes now to simplify

OpenStudy (amistre64):

add 0 :)

OpenStudy (anonymous):

can't be simplify further: http://www.wolframalpha.com/input/?i=(x-x^(1%2F4))%2F(x-1)&t=crmtb01

OpenStudy (anonymous):

to find lim >1

OpenStudy (amistre64):

\[1-\cfrac{1}{4}x^{-3/4}\] id say its 3/4

OpenStudy (anonymous):

you are correct thx

OpenStudy (amistre64):

L'Hopitals rule for indeterminate forms

myininaya (myininaya):

do you want a non L'hospital way?

OpenStudy (anonymous):

kk

myininaya (myininaya):

\[Let u=x^4\] as x->1, u->1 \[u=x^\frac{1}{4} =>u^4=x\] so we have \[\lim_{u \rightarrow 1}\frac{u^4-u}{u^4-1}=\lim_{x \rightarrow 1}\frac{u(u^3-1)}{(u^2-1)(u^2+1)}\] \[=\lim_{u \rightarrow 1}\frac{u(u-1)(u^2+u+1)}{(u-1)(u+1)(u^2+1)}=\lim_{u \rightarrow 1}\frac{u(u^2+u+1)}{(u+1)(u^2+1)}\]

myininaya (myininaya):

\[=\frac{1(1^2+1+1)}{(1+1)(1^2+1)}=\frac{1(3)}{2(2)}=\frac{3}{4}\]

OpenStudy (anonymous):

thx

myininaya (myininaya):

np